# Joint Learning of Linear Dynamical Systems under Smoothness Constraints

### v1

@@ 3.2 Subspace-constrained LS
$\frac{1}{m}\sum_{l=1}^{m}\big{\|}{\widehat{A}_{l}-A_{l}^{*}}\big{\|}_{F}^{2}\leq C_{1}\frac{S_{m}}{m\lambda_{m-\tau}}\left[1+\frac{\gamma_{1}^{2}(T,m,\delta)}{T^{2}}\right]\mathbf{1}_{\left\{{\tau\leq m-1}\right\}}+C_{2}\frac{R^{2}d^{2}\tau}{Tm}\log\left(\frac{\theta\tau\gamma_{3}(m,T)}{mT}+\frac{1}{\delta}\right).$ (3.15)
@@
    $\frac{1}{m}\sum_{l=1}^{m}\big{\|}{\widehat{A}_{l}-A_{l}^{*}}\big{\|}_{F}^{2}\leq C_{1}\frac{mS_{m}}{(\tau^{*})^{2}}\left[1+\frac{\gamma_{1}^{2}(T,m,\delta)}{T^{2}}\right]\mathbf{1}_{\left\{{\tau^{*}\leq m-1}\right\}}+C_{2}\frac{R^{2}d^{2}\tau^{*}}{Tm}\log\left(\frac{\theta\tau^{*}\gamma_{3}(m,T)}{mT}+\frac{1}{\delta}\right).$ (3.19)
@@
    $\tau=\left\lfloor\left(\frac{mT}{C_{3}^{\prime}dR^{4}\log\Big{(}\frac{dmT^{\alpha}\kappa^{2}_{\max}}{\delta}\Big{)}}\right)^{1/3}\right\rfloor$
@@
    $\frac{1}{m}\sum_{l=1}^{m}\big{\|}{\widehat{A}_{l}-A_{l}^{*}}\big{\|}_{F}^{2}\leq C_{4}R^{20/3}d^{11/3}\kappa^{4}_{\max}T^{2\alpha-\frac{2}{3}}\log^{2}(\frac{m}{\delta})\Big{(}m^{1/3}S_{m}+m^{-1/3}\Big{)}\log^{2/3}\Big{(}\frac{dmT^{\alpha}\kappa^{2}_{\max}}{\delta}\Big{)}.$
@@
    The second part of the above corollary is written for the regime $S_{m}=\Omega(1/m)$, but it is not hard to derive error rates for the setting where $S_{m}=O(1/m)$. Notice that in order to ensure $\frac{1}{m}\sum_{l=1}^{m}\big{\|}{\widehat{A}_{l}-A_{l}^{*}}\big{\|}_{F}^{2}=o(1)$ as $m\rightarrow\infty$, we require $S_{m}=o(m^{-1/3})$, which, as alluded to earlier, is a consequence of the condition (3.14). To illustrate this smoothness value, recall Example 1 – in that setting, this would mean that each $A^{*}_{l,i,j}$ is formed by sampling a $(L,\beta)$-Hölder function with $\beta\in(2/3,1]$.
@@
    $\frac{1}{m}\sum_{l=1}^{m}\big{\|}{\widehat{A}_{l}-A_{l}^{*}}\big{\|}_{F}^{2}=O\Big{(}4^{\alpha}d^{11/3}\log^{8/3}\Big{(}\frac{dm2^{\alpha}}{\delta}\Big{)}(m^{1/3}S_{m}+m^{-1/3})\Big{)}.$
    For instance, if $S_{m}=O(m^{-2/3})$, this implies that the MSE is $O(4^{\alpha}d^{11/3}\log^{8/3}(\frac{dm2^{\alpha}}{\delta})m^{-1/3})$. If $\alpha\asymp d$, then $m$ is required to grow exponentially with $d$ for controlling the MSE which is again an artefact of the analysis. It arises due to the $\gamma_{1}(T,m,\delta)$ term appearing outside the logarithm in the first term in (3.15).
@@ 4.2 Proof of Corollary 1
Next, using the fact $\sin x\in[\frac{2x}{\pi},x]$ for $x\in[0,\pi/2]$, this implies $\lambda_{l}\in[c_{1}\frac{(m-l)^{2}}{m^{2}},c_{2}\frac{(m-l)^{2}}{m^{2}}]$ for each $l$. Hence we can bound $\sum_{l=1}^{m-1}\frac{1}{4\lambda\lambda_{l}+T}$ bound as
@@ 4.3.2 Proof for star graph
Recall that for a star graph, the eigenvalues of $L$ are given by Brouwer and Haemers (2012)
$\lambda_{m}=0,\ \lambda_{1}=m,\ \text{ and }\lambda_{l}=1\text{ for }l=2,\dots,m-1.$
This leads to the bound
$\big{\|}{(L\otimes I_{d^{2}})a^{*}}\big{\|}_{2}^{2}=\sum_{l=1}^{m}\sum_{i=1}^{d^{2}}\lambda_{l}^{2}\Big{\langle}a^{*},v_{l}\otimes e_{i}\Big{\rangle}^{2}=m^{2}\sum_{i=1}^{d^{2}}\Big{\langle}a^{*},v_{1}\otimes e_{i}\Big{\rangle}^{2}+\sum_{l=2}^{m-1}\sum_{i=1}^{d^{2}}\Big{\langle}a^{*},v_{l}\otimes e_{i}\Big{\rangle}^{2}\leq\tilde{S}_{m},$
where the last bound follows from Assumption 2. It is also easy to verify that $\sum_{l=1}^{m-1}\frac{1}{4\lambda\lambda_{l}+T}\leq\frac{2m}{\lambda+T}$. Using these bounds in (1) and the fact $\gamma_{1}(T,m,\delta)\geq 1$, we obtain
@@ 5.1 Proof of Theorem 2
Before proceeding, it will be useful to denote $V_{\tau}\in\mathbb{R}^{md^{2}\times\tau d^{2}}$ to be an orthonormal-basis matrix for $\text{span}(\widetilde{P}^{(\tau)})$. In fact given the definition of $\widetilde{P}^{(\tau)}$, we can explicitly write $V_{\tau}$ as
@@
Let $G$ be $\theta$-delocalized. Denote $\mathcal{E}_{3}$ to be the event
@@ 5.2.2 Proof of Part $2$
$\gamma_{i}(m,T)\lesssim T\sum_{l=1}^{m}dT^{\alpha_{l}}\kappa^{2}(P_{l})\leq dmT^{\alpha+1}\kappa^{2}_{\max},\quad\text{for $i=1,2$}.$ (5.12)
@@
The regime for $S_{m}$ in the statement follows from the condition $\bar{\tau}^{*}\leq\Big{(}\frac{2m^{2}S_{m}T}{R^{2}d^{2}}\Big{)}^{1/3}$ with $\bar{\tau}^{*}$ as in (5.14), which implicitly implies $\Big{(}\frac{2m^{2}S_{m}T}{R^{2}d^{2}}\Big{)}^{1/3}\geq 1$ (since $\bar{\tau}^{*}\geq 1$). Hence in this smoothness regime, $\tilde{\tau}^{*}=\bar{\tau}^{*}$. Plugging $\tau^{*}=\tau^{*}=\lfloor\tilde{\tau}^{*}\rfloor$ in (1), and using the bounds in (5.12), (5.13), we obtain the stated error bound after some minor simplifications.
@@ A.1 Proof of Proposition 1
Denote $e_{1},\dots,e_{m}\in\mathbb{R}^{m}$ to be the canonical basis. Using the definition of $Q$, we can write
@@
Denote $(\mathcal{F}_{t})_{t=1}^{\infty}$ to be a filtration where $\mathcal{F}_{t}:=\sigma(\eta_{l,1},\dots,\eta_{l,t})_{l=1}^{m}$.
@@
meaning that $(M_{t}^{\mu})_{t=1}^{\infty}$ is a supermartingale, and also $\mathbb{E}[M_{t}^{\mu}]\leq 1$. The arguments for showing that $M_{\overline{\tau}}^{\mu}$ is well-defined, and $\mathbb{E}[M_{\overline{\tau}}^{\mu}]\leq 1$ is identical to (Abbasi-yadkori et al., 2011, Lemma 9), and hence omitted. ∎
@@ Appendix C Useful tools
AtF2

### v2

@@ 3.2 Subspace-constrained LS
$\frac{1}{m}\sum_{l=1}^{m}\big{\|}{\widehat{A}_{l}-A_{l}^{*}}\big{\|}_{F}^{2}\leq C_{1}\frac{S_{m}}{m\lambda_{m-\tau}}\left[1+\frac{\gamma_{1}^{2}(T,m,\delta)}{T^{2}}\right]\mathbf{1}_{\left\{{\tau\leq m-1}\right\}}+C_{2}\frac{R^{2}d^{2}\tau}{Tm}\log\left(\frac{\theta\tau\gamma_{3}(m,T)}{{\color[rgb]{0,0,0}\delta}mT}+\frac{1}{\delta}\right).$ (3.15)
@@
    $\frac{1}{m}\sum_{l=1}^{m}\big{\|}{\widehat{A}_{l}-A_{l}^{*}}\big{\|}_{F}^{2}\leq C_{1}\frac{mS_{m}}{(\tau^{*})^{2}}\left[1+\frac{\gamma_{1}^{2}(T,m,\delta)}{T^{2}}\right]\mathbf{1}_{\left\{{\tau^{*}\leq m-1}\right\}}+C_{2}\frac{R^{2}d^{2}\tau^{*}}{Tm}\log\left(\frac{\tau^{*}\gamma_{3}(m,T)}{{\color[rgb]{0,0,0}\delta}mT}+\frac{1}{\delta}\right).$ (3.19)
@@
    ${\color[rgb]{0,0,0}\tau^{*}}=\left\lfloor\left(\frac{mT}{C_{3}^{\prime}dR^{4}\log\Big{(}\frac{dmT^{\alpha}\kappa^{2}_{\max}}{\delta}\Big{)}}\right)^{1/3}\right\rfloor$
@@
    $\frac{1}{m}\sum_{l=1}^{m}\big{\|}{\widehat{A}_{l}-A_{l}^{*}}\big{\|}_{F}^{2}\leq C_{4}R^{20/3}{\color[rgb]{0,0,0}d^{8/3}}\kappa^{4}_{\max}T^{2\alpha-\frac{2}{3}}\log^{2}(\frac{m}{\delta})\Big{(}m^{1/3}S_{m}+{\color[rgb]{0,0,0}m^{-2/3}}\Big{)}\log^{2/3}\Big{(}\frac{dmT^{\alpha}\kappa^{2}_{\max}}{\delta}\Big{)}.$ (3.21)
@@
    The bound in (3.21) increases with $T$, which is essentially due to the $\gamma_{1}(T,m,\delta)$ term in (3.19). This is an artefact of the analysis, and it is unclear how to improve this for the case of marginally stable matrices. But Remark 3 applies here as well – in case $\big{\|}{A^{*}_{l}}\big{\|}_{2}<1$ for each $l$, we can simply set $\alpha=0$ and $\kappa^{2}_{\max}=(1-\max_{l}\big{\|}{A^{*}_{l}}\big{\|}_{2}^{2})^{-1}$ in Corollary 3. Then, the bound on the MSE will be a decreasing function of $T$.
@@
    The second part of the above corollary is written for the regime $S_{m}=\Omega(1/m)$, but it is not hard to derive error rates for the setting where $S_{m}=O(1/m)$. Notice that in order to ensure $\frac{1}{m}\sum_{l=1}^{m}\big{\|}{\widehat{A}_{l}-A_{l}^{*}}\big{\|}_{F}^{2}=o(1)$ as $m\rightarrow\infty$, we require $S_{m}=o(m^{-1/3})$, which, as alluded to earlier, is a consequence of the condition (3.14). To illustrate this smoothness value, recall Example 1 – in that setting, this would mean that each $A^{*}_{l,i,j}$ is formed by sampling a $(M,\beta)$-Hölder function with $\beta\in(2/3,1]$.
  3. 3.
@@
    $\frac{1}{m}\sum_{l=1}^{m}\big{\|}{\widehat{A}_{l}-A_{l}^{*}}\big{\|}_{F}^{2}=O\Big{(}4^{\alpha}d^{11/3}\log^{8/3}\Big{(}\frac{dm2^{\alpha}}{\delta}\Big{)}(m^{1/3}S_{m}+{\color[rgb]{0,0,0}m^{-2/3}})\Big{)}.$
    For instance, if $S_{m}=O(m^{-2/3})$, this implies that the MSE is $O(4^{\alpha}d^{11/3}\log^{8/3}(\frac{dm2^{\alpha}}{\delta})m^{-1/3})$. If $\alpha\asymp d$, then $m$ is required to grow exponentially with $d$ for controlling the MSE.
@@ 4.2 Proof of Corollary 1
Next, using the fact $\sin x\in[\frac{2x}{\pi},x]$ for $x\in[0,\pi/2]$, this implies $\lambda_{l}\in[c_{1}\frac{(m-l)^{2}}{m^{2}},c_{2}\frac{(m-l)^{2}}{m^{2}}]$ for each $l$. Hence we can bound $\sum_{l=1}^{m-1}\frac{1}{4\lambda\lambda_{l}+T}$ as
@@ 4.3.2 Proof for star graph
Recall that for a star graph, the eigenvalues of $L$ are given by Spielman (2025)
$\lambda_{m}=0,\ \lambda_{1}=m,\ \text{ and }\lambda_{l}=1\text{ for }l=2,\dots,m-1$
and also $v_{1}=[m(m-1)]^{-1/2}(-(m-1),1,\cdots,1)^{\top}$. This leads to the bound
$\big{\|}{(L\otimes I_{d^{2}})a^{*}}\big{\|}_{2}^{2}$ $=\sum_{l=1}^{m}\sum_{i=1}^{d^{2}}\lambda_{l}^{2}\Big{\langle}a^{*},v_{l}\otimes e_{i}\Big{\rangle}^{2}$
$=m^{2}\sum_{i=1}^{d^{2}}\Big{\langle}a^{*},v_{1}\otimes e_{i}\Big{\rangle}^{2}+\sum_{l=2}^{m-1}\sum_{i=1}^{d^{2}}\Big{\langle}a^{*},v_{l}\otimes e_{i}\Big{\rangle}^{2}$
$\leq{\color[rgb]{0,0,0}(m^{2}-m)\sum_{i=1}^{d^{2}}\Big{\langle}a^{*},v_{1}\otimes e_{i}\Big{\rangle}^{2}+S_{m}}$
$\leq{\color[rgb]{0,0,0}m(m-1)(a^{*})^{\top}[(v_{1}\otimes I_{d^{2}})(v_{1}^{\top}\otimes I_{d^{2}})]a^{*}+S_{m}}$
$\leq{\color[rgb]{0,0,0}\tilde{S}_{m}},$
where the last equality follows from the expression for $v_{1}$ with some minor simplification. It is also easy to verify that $\sum_{l=1}^{m-1}\frac{1}{4\lambda\lambda_{l}+T}\leq\frac{2m}{\lambda+T}$. Using these bounds in (1) and the fact $\gamma_{1}(T,m,\delta)\geq 1$, we obtain
@@ 5.1 Proof of Theorem 2
Before proceeding, it will be useful to denote an orthonormal-basis matrix for $\text{span}(\widetilde{P}^{(\tau)})$ by $V_{\tau}\in\mathbb{R}^{md^{2}\times\tau d^{2}}$. In fact given the definition of $\widetilde{P}^{(\tau)}$, we can explicitly write $V_{\tau}$ as
@@
Let $G$ be $\theta$-delocalized and denote the event
@@ 5.2.2 Proof of Part $2$
$\gamma_{i}(m,T)\lesssim T\sum_{l=1}^{m}dT^{\alpha_{l}}\kappa^{2}(P_{l})\leq dmT^{\alpha+1}\kappa^{2}_{\max},\quad\text{for {\color[rgb]{0,0,0}$i=2,3$}}.$ (5.12)
@@
The regime for $S_{m}$ in the statement follows from the condition $\bar{\tau}^{*}\leq\Big{(}\frac{2m^{2}S_{m}T}{R^{2}d^{2}}\Big{)}^{1/3}$ with $\bar{\tau}^{*}$ as in (5.14), which implicitly implies $\Big{(}\frac{2m^{2}S_{m}T}{R^{2}d^{2}}\Big{)}^{1/3}\geq 1$ (since $\bar{\tau}^{*}\geq 1$). Hence in this smoothness regime, $\tilde{\tau}^{*}=\bar{\tau}^{*}$. Plugging $\tau^{*}=\lfloor\tilde{\tau}^{*}\rfloor$ in (1), and using the bounds in (5.12), (5.13) and (4.6), we obtain the stated error bound after some minor simplifications.
@@ A.1 Proof of Proposition 1
Denote the canonical basis of $\mathbb{R}^{m}$ by $e_{1},\dots,e_{m}$. Using the definition of $Q$, we can write
@@
Denoting $\mathcal{F}_{t}:=\sigma(\eta_{l,1},\dots,\eta_{l,t})_{l=1}^{m}$, we obtain a filtration $(\mathcal{F}_{t})_{t=1}^{\infty}$.
@@
meaning that $(M_{t}^{\mu})_{t=1}^{\infty}$ is a supermartingale, and also $\mathbb{E}[M_{t}^{\mu}]\leq 1$. The arguments for showing that $M_{\overline{\tau}}^{\mu}$ is well-defined, and $\mathbb{E}[M_{\overline{\tau}}^{\mu}]\leq 1$ are identical to (Abbasi-yadkori et al., 2011, Lemma 9), and hence omitted. ∎
@@ Appendix C Useful tools
$\big{\|}{A^{t}}\big{\|}_{F}^{2}=\big{\|}{P^{-1}\Lambda^{t}P}\big{\|}_{F}^{2}\leq\big{\|}{P^{-1}}\big{\|}_{2}^{2}\big{\|}{P}\big{\|}_{2}^{2}\big{\|}{\Lambda^{t}}\big{\|}_{F}^{2},$
which in turn implies
$\operatorname{Tr}(\Gamma_{T}(A)-I)=\sum_{t=1}^{T}\operatorname{Tr}(A^{t}(A^{t})^{\top})\leq\big{\|}{P^{-1}}\big{\|}_{2}^{2}\big{\|}{P}\big{\|}_{2}^{2}\sum_{t=1}^{T}\operatorname{Tr}(\Lambda^{t}(\Lambda^{t})^{H}).$
Here ${(\Lambda^{t})^{H}}$ denotes the Hermitian transpose of $\Lambda^{t}$ as $\Lambda$ may have complex-value entries. Since
$\Lambda^{t}(\Lambda^{t})^{H}=\begin{pmatrix}\sum_{t=1}^{T}J^{t}_{k_{1}}(\lambda_{1})(J^{t}_{k_{1}}(\lambda_{1}))^{H}&&&\\ &\ddots&&\\ &&\ddots&\\ &&&\sum_{t=1}^{T}J^{t}_{k_{s}}(\lambda_{s})(J^{t}_{k_{s}}(\lambda_{s}))^{H}\end{pmatrix},$
it will suffice to bound $\operatorname{Tr}(\sum_{t=1}^{T}J^{t}_{k_{1}}(\lambda_{1})(J^{t}_{k_{1}}(\lambda_{1}))^{H})$. Now as mentioned in (Sarkar and Rakhlin, 2019, Proposition 7.5), there is a $k_{1}\times k_{1}$ unitary diagonal matrix $D$ such that $D(\lambda_{1}I_{k_{1}}+N_{k_{1}})(\lambda_{1}I_{k_{1}}+N_{k_{1}})^{H}D^{H}$ is a real matrix, and in particular,
$D(\lambda_{1}I_{k_{1}}+N_{k_{1}})(\lambda_{1}I_{k_{1}}+N_{k_{1}})^{H}D^{H}=(\left|{\lambda_{1}}\right|I_{k_{1}}+N_{k_{1}})(\left|{\lambda_{1}}\right|I_{k_{1}}+N_{k_{1}})^{H}.$
Thus we obtain
$\operatorname{Tr}\left(\sum_{t=1}^{T}J^{t}_{k_{1}}(\lambda_{1})(J^{t}_{k_{1}}(\lambda_{1}))^{H}\right)$ $=\operatorname{Tr}\left(D\left[\sum_{t=1}^{T}J^{t}_{k_{1}}(\lambda_{1})(J^{t}_{k_{1}}(\lambda_{1}))^{H}\right]D^{H}\right)$
$=\operatorname{Tr}\left(\underbrace{\sum_{t=1}^{T}(\left|{\lambda_{1}}\right|I_{k_{1}}+N_{k_{1}})(\left|{\lambda_{1}}\right|I_{k_{1}}+N_{k_{1}})^{H}}_{=:B}\right)$
$=\sum_{l=1}^{k_{1}}B_{ll}.$
If $\left|{\lambda_{1}}\right|\leq 1+\frac{c}{T}$, then eq. (34) in the proof of (Sarkar and Rakhlin, 2019, Proposition 7.5) implies
$B_{ll}\leq e^{2c}T\sum_{j=0}^{k_{1}-l}T^{2j}=e^{2c}T\left(\frac{T^{2(k_{1}-l+1)}-1}{T^{2}-1}\right)\lesssim e^{2c}T^{2(k_{1}-l)+1},$ (C.1)
if $T\geq 2$. This in turn implies
$\sum_{l=1}^{k_{1}}B_{ll}\lesssim e^{2c}\sum_{l=1}^{k_{1}}T^{2(k_{1}-l)+1}\lesssim e^{2c}T^{2k_{1}-1}$
which readily leads to the bound in the first part of the proposition.
When $T=1$, notice from the first inequality in (C.1) that $B_{ll}\leq e^{2c}(k_{1}-l+1)$, and so,
$\sum_{l=1}^{k_{1}}B_{ll}\leq e^{2c}\sum_{l=1}^{k_{1}}(k_{1}-l+1)=\sum_{i=1}^{k_{1}}i\leq e^{2c}k_{1}^{2}.$
This leads to the stated bound in the second part of the proposition. ∎